OpenStudy (anonymous):

f(x)=x^2/1+0.25x^2 a) find the domain of f b) find any horizontal or vertical asymptotes

OpenStudy (anonymous):

Show the function as I think it was intended, as opposed to what you typed in:\[f(x)=x^2/(1+0.25x^2)=x^2/(1+x^2/4)\] a) The domain of f is all the numbers that can be successfully put in for x. Usually, we delete those numbers that create a zero in the denominator, but there is no such real number, so\[dom(f)=\mathbb{R}=(-\infty,\infty)\] b)Vertical asymptotes occur when the denominator goes to zero. As discussed above, there is no such value of x, so no vertical asymptotes. Horizontal asymptotes happen as x gets really positive or negative, and the degree of the numerator and denominator are the same. That is not the case here, so there are no horizontal asymptotes, either.